# via Continuized Discrete Diffusion

### v1

@@ 8.1 Architecture and Pipeline
During inference, we begin from an initial noisy action sequence and perform a multi-step iterative denoising procedure. At the first iteration, the observation $o_{t}$ is encoded and processed by the model to produce the initial key–value cache $\mathit{KV}(o_{t})$, which stores the key–value pairs required for cross-attention. This cache is then reused across all subsequent denoising steps, so that only the action tokens need to be recomputed.
@@
At each iteration $i\in\{1,\ldots,N\}$, the current noisy action sequence $\tilde{A}_{t}^{\tau_{i}}$ is passed into the model together with the fixed cache $\mathit{KV}(o_{t})$. The model produces logits $v_{\theta}(\tilde{A}_{t}^{\tau_{i}},\mathit{KV}(o_{t}))$, which define a categorical distribution over actions:
@@
The categorical distribution is decoded into discrete token indices, each of which is converted into a one-hot representation:
$\mathrm{onehot}(k)\in\{0,1\}^{B},\qquad\mathrm{onehot}(k)_{j}=\begin{cases}1,&j=k,\\ 0,&j\neq k,\end{cases}$ (11)
with $B$ denoting the number of discretization bins. The resulting one-hot tokens form the intermediate representation $\hat{A}_{t}^{(i)}$. To generate the next noisy input, we reapply the forward noise process:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\hat{A}_{t}^{(i)}+(1-\tau_{i+1})\,\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I).$ (12)
After $N$ iterations, the final predicted token indices $\{k_{h}\}_{h=1}^{H}$ are deterministically mapped back into continuous actions by inverting the discretization:
$a_{t,h}=\mathrm{Detokenize}(k_{h}),\qquad h=1,\ldots,H.$ (13)
Collecting across the horizon yields the reconstructed action chunk:
@@
8.2 Spherical Perspective Generalization
@@
Spherical Warping.
Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and camera intrinsics $(f_{x},f_{y},c_{x},c_{y})$, each pixel $(u,v)$ is first unprojected to a 3D point at fixed depth $d_{\mathrm{center}}$:
$X=\frac{(u-c_{x})\,d_{\mathrm{center}}}{f_{x}},\qquad Y=\frac{(v-c_{y})\,d_{\mathrm{center}}}{f_{y}},\qquad Z=d_{\mathrm{center}}.$ (15)
@@
$R=R_{x}(\Delta\phi)\,R_{y}(\Delta\theta),\qquad\begin{bmatrix}X^{\prime}\\ Y^{\prime}\\ Z^{\prime}\end{bmatrix}=R\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$. (16)
@@
$u^{\prime}=\frac{X^{\prime}f_{x}}{Z^{\prime}}+c_{x},\qquad v^{\prime}=\frac{Y^{\prime}f_{y}}{Z^{\prime}}+c_{y},$ (17)
@@
Relative Spherical Embedding.
To explicitly encode the relative perturbation of each camera, we define a 3D offset vector $\delta=[d,\;\theta,\;\phi]$, where $(d,\theta,\phi)$ denote radial, horizontal, and vertical displacements. For the base camera, we use $\delta_{\mathrm{orig}}=[-d,-\theta,-\phi]$, and for the warped camera, $\delta_{\mathrm{warp}}=[d,\theta,\phi]$. These offsets are mapped into the token embedding space through a learnable projection function $f_{\mathrm{proj}}$:
@@
9 Detailed Ablation Analysis
Effect of Discretization Bins.
A key advantage of our model is its ability to flexibly discretize continuous control signals into an arbitrary number of bins, enabling a fine-grained trade-off between numerical precision and representation compactness. To evaluate this, we systematically vary the discretization resolution from 128 to 8192 bins (see Tab. 4). At low resolutions (e.g., 128 bins), the model fails to learn meaningful trajectories due to the inability of coarse quantization to accurately encode joint values. This results in erratic and unstable robot motions. As the bin count increases, control accuracy improves, with performance saturating around 2048 bins. At this point, the discretization error becomes negligible compared to diffusion noise, offering sufficient precision for smooth and stable control. However, further increasing the resolution to 4096 or 8192 bins yields only marginal gains or even slight degradation in performance. This is attributed to the larger token vocabulary and increased embedding noise, which complicate optimization. Unlike conventional autoregressive VLA models limited to 256 discrete bins, our method supports significantly higher-resolution action representations, enabling superior control fidelity and precision.
@@ 9 Detailed Ablation Analysis
Effect of Action Dimension.
Our model supports flexible action dimensionality when discretizing and embedding continuous control vectors. To assess the influence of this flexibility, we vary the number of action dimensions $d\in\{8,16,32\}$ during training and evaluation (see Tab. 5). We observe that setting $d$ to match or slightly exceed the maximum action dimensionality in the dataset (e.g., d=8d{=}8 in LIBERO) yields the highest average success rate. When $d$ is smaller than the true action dimension, the model cannot capture all control channels, leading to missing or corrupted actions during decoding. Conversely, using a value of $d$ significantly larger than necessary introduces numerous blank or zero-padded tokens into the sequence. These redundant tokens inject unnecessary noise into the diffusion process, degrading the sharpness and precision of predicted trajectories. Therefore, for optimal performance, $d$ should be chosen to tightly align with the dataset’s action space dimensionality, ensuring both representational completeness and robustness against embedding-induced noise.
@@
Effect of One-hot Smooth Factor.
Prior to injecting noise into the one-hot action representation, we apply a one-hot smooth factor to scale the vector. This smoothing operation attenuates the sharp peaks of the one-hot distribution, thereby reducing the dominance of high-confidence logits and enabling smoother transitions within the action space during diffusion. As shown in Tab. 6, lower values of the smooth factor consistently improve success rates across all LIBERO subsets. Stronger attenuation encourages the model to better explore neighboring discrete actions and prevents overfitting to rigid one-hot boundaries. Conversely, higher smooth factor values retain excessively sharp logits, which hinder denoising stability and degrade generalization performance. These findings indicate that moderately suppressing one-hot amplitude prior to noise injection enhances the robustness of policy learning under diffusion-based action modeling.
@@
Effect of Action Chunk.
We adopt an action chunking strategy, wherein the policy outputs a short sequence of future actions conditioned on the current observation. This approach has shown particular effectiveness in real-world imitation learning and VLA models (e.g., [chi2023diffusion, liu2024rdt, black2024pi_0]). By predicting multiple future actions simultaneously, the policy captures implicit temporal context, thereby mitigating non-Markovian ambiguities such as pauses, delayed responses, or lingering motions—without requiring additional historical inputs that may introduce causal confusion.
@@
Effect of Data Normalization.
We compare two normalization techniques for processing continuous action vectors within our model. Both methods apply element-wise normalization based on statistical parameters (e.g., mean and standard deviation) computed independently for each action dimension across the entire dataset.
@@
10 Hardware Details
We conducted real-world experiments using a Franka Emika Research 3 robotic arm, equipped with two Intel RealSense D435i cameras. One camera was mounted on the robot’s wrist, providing a dynamic view that changes with the motion of the robotic arm. The second camera was placed laterally in the workspace to capture a stable third-person perspective. We utilized Gello [wu2024gello] to control the Franka arm and collect demonstration data.
@@ 10 Hardware Details
11 More Details on LIBERO
To evaluate the generalization behavior of our model under diverse distribution shifts, we adopt LIBERO [liu2023libero] as a comprehensive benchmark for robot manipulation. LIBERO provides 130 language-conditioned tasks grouped into four procedurally generated suites (spatial, object, goal, and libero-100), each targeting a different dimension of knowledge transfer, including spatial reasoning, object concepts, behavioral goals, and mixed declarative–procedural knowledge. These tasks are created through a scalable generation pipeline that samples human-activity-inspired templates, constructs scene layouts, and specifies PDDL-based goal predicates, resulting in highly diverse manipulation scenarios with controllable variations in object types, spatial layouts, and goal semantics.
Following the common practice in prior work (use libero-10 instead of libero-100) and consistent with the configurations adopted by most existing models [black2024pi_0, kim2024openvla, pertsch2025fast], we train a single unified policy across all LIBERO tasks, instead of training task-specific models. All training is performed using a single RTX PRO 6000 GPU over 30,000 steps. Each task provides 50 high-quality teleoperated demonstrations, enabling sample-efficient behavioral cloning and ensuring that performance differences primarily reflect the model’s ability to transfer and retain knowledge across tasks.
@@ 11 More Details on LIBERO
Diffusion Policy [chi2023diffusion] 78.3 92.5 68.3 50.5 72.4
MDT [reuss2024multimodal] 78.5 87.5 73.5 64.8 76.1
OpenVLA [kim2024openvla] 84.7 88.4 79.2 53.7 76.5
Octo [mees2024octo] 78.9 85.7 84.6 51.1 75.1
Dita [hou2025dita] 84.2 96.3 85.4 63.8 82.4
TraceVLA [zheng2024tracevla] 84.6 85.2 75.1 54.1 74.8
SpatialVLA [qu2025spatialvla] 88.2 89.9 78.6 55.5 78.1
$\pi_{0}$ FAST [pertsch2025fast] 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ [black2024pi_0] 96.8 98.8 95.8 85.2 94.2
$\pi_{0.5}$ [intelligence2025pi05] 95.4 98.4 97.2 89.6 95.2
[…]
12 More Details on ManiSkill
To further assess the generalization capability and fine-grained manipulation performance of our model across diverse operational scenarios, we additionally conduct experiments on ManiSkill [tao2024maniskill3]. ManiSkill is an advanced GPU–parallelized simulation and benchmarking framework for embodied AI, offering one of the most comprehensive suites of manipulation tasks to date. It covers 12 distinct categories, ranging from tabletop manipulation and dexterous hand control to room-scale mobile manipulation, bimanual coordination, and high-fidelity digital twin environments. Its highly efficient GPU-based simulation and rendering pipeline enables large-scale visual policy training on a single GPU with minimal memory overhead, making it a suitable platform for evaluating both generalization and visual robustness.
Following the experimental design used in RDT [liu2024rdt], we leverage the official demonstration dataset (5 tasks with a total of 5,000 trajectories) and retrain all baseline models under identical data conditions to ensure fair comparison. All training is performed on a single RTX PRO 6000 GPU for 30,000 steps, using the same evaluation protocol across models. This setup allows us to rigorously compare their representational capacity, generalization behavior, and fine-grained manipulation performance within a high-fidelity simulation environment.
@@ 12 More Details on ManiSkill
Diffusion Policy [chi2023diffusion] 0.0 40.0 80.0 0.0 88.0 30.2
OpenVLA [kim2024openvla] 0.0 8.0 80.0 0.0 8.0 4.8
Octo [mees2024octo] 0.0 0.0 0.0 0.0 0.0 0.0
RDT [liu2024rdt] 13.2 77.2 74.0 1.2 100.0 53.6
$\pi_{0}$ FAST [pertsch2025fast] 0.0 80.0 52.0 0.0 92.0 44.8
$\pi_{0}$ [black2024pi_0] 4.0 60.0 48.0 0.0 100.0 42.4
$\pi_{0.5}$ [intelligence2025pi05] 8.0 56.0 56.0 4.0 92.0 43.2
[…]
13 More Details on VLABench
To further assess the robustness, semantic understanding, and generalization capability of our model across diverse language-conditioned manipulation scenarios, we additionally evaluate it on VLABench [zhang2024vlabench]. VLABench offers a broad and heterogeneous collection of tasks that span various object categories, spatial arrangements, and visually complex environments. Each task is paired with natural language instructions that require accurate semantic grounding, attribute-level understanding, and precise alignment between linguistic cues and visual observations. The benchmark’s extensive variability in object layouts, appearance conditions, and linguistic expressions makes it a rigorous platform for evaluating multimodal reasoning and visuomotor robustness.

### v2

@@ A.1 Architecture and Pipeline
During inference, we initialize the action sequence from Gaussian noise and perform a multi-step iterative refinement procedure. At the first iteration, the observation $o_{t}$ is encoded once by the model to construct a key–value cache $\mathit{KV}(o_{t})$ for cross-attention. This cache is reused across all subsequent refinement steps, such that only the action tokens are recomputed.
@@
We discretize a continuous refinement variable $\tau\in[1,0]$ into $N$ (with $N=10$ in our experiments) steps using a uniform schedule, i.e., $\tau_{i}=1-\frac{i-1}{N}$ for $i=1,\ldots,N$. At each iteration $i$, the current noisy action representation $\tilde{A}_{t}^{\tau_{i}}$ is passed into the model together with the fixed cache $\mathit{KV}(o_{t})$. The model produces logits $v_{\theta}(\tilde{A}_{t}^{\tau_{i}},\mathit{KV}(o_{t}))$, which define a categorical distribution over discrete action tokens:
@@
The predicted categorical distribution is decoded into discrete token indices via $\arg\max$ decoding, and each token is converted into a one-hot vector in a $B$-dimensional discrete action space:
$\mathrm{onehot}(k)\in\{0,1\}^{B},\quad\mathrm{onehot}(k)_{j}=\begin{cases}1,&j=k,\\ 0,&j\neq k.\end{cases}$ (11)
The resulting one-hot vectors form the intermediate action representation $\hat{A}_{t}^{(i)}$.
To obtain the input for the next refinement step, we apply an implicit noise annealing scheme controlled by $\tau_{i+1}$:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\varepsilon+(1-\tau_{i+1})\,\hat{A}_{t}^{(i)},\quad\varepsilon\sim\mathcal{N}(0,I),$ (12)
where $\varepsilon$ is fixed-variance isotropic Gaussian noise. As $\tau$ decreases monotonically, the signal-to-noise ratio increases, yielding progressively more deterministic action representations without introducing an explicit variance schedule.
After $N$ refinement steps, the final predicted token indices $\{k_{h}\}_{h=1}^{H}$ are deterministically mapped back into continuous actions by inverting the discretization:
$a_{t,h}=\mathrm{Detokenize}(k_{h}),\quad h=1,\ldots,H.$ (13)
In practice, following (Black et al., 2024; Intelligence et al., 2025), we precompute per-dimension action statistics and normalize all action values into the $[0,1]$ interval prior to discretization. This preprocessing enables efficient quantile-based discretization, where actions can be easily encoded into discrete tokens and inverse-mapped back to the original continuous action space through the corresponding rescaling operation. Collecting across the horizon yields the reconstructed action chunk:
@@
A.2 Spherical Perspective Generalization
@@
Spherical Warping. Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and camera intrinsics $(f_{x},f_{y},c_{x},c_{y})$, each pixel $(u,v)$ is first unprojected to a 3D point at fixed depth $d_{\mathrm{center}}$:
$X=\frac{(u-c_{x})\,d_{\mathrm{center}}}{f_{x}},\quad Y=\frac{(v-c_{y})\,d_{\mathrm{center}}}{f_{y}},\quad Z=d_{\mathrm{center}}.$ (15)
@@
$R=R_{x}(\Delta\phi)\,R_{y}(\Delta\theta),\quad\begin{bmatrix}X^{\prime}\\ Y^{\prime}\\ Z^{\prime}\end{bmatrix}=R\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$. (16)
@@
$u^{\prime}=\frac{X^{\prime}f_{x}}{Z^{\prime}}+c_{x},\quad v^{\prime}=\frac{Y^{\prime}f_{y}}{Z^{\prime}}+c_{y},$ (17)
@@
Relative Spherical Embedding. To explicitly encode the relative perturbation of each camera, we define a 3D offset vector $\delta=[d,\;\theta,\;\phi]$, where $(d,\theta,\phi)$ denote radial, horizontal, and vertical displacements. For the base camera, we use $\delta_{\mathrm{orig}}=[-d,-\theta,-\phi]$, and for the warped camera, $\delta_{\mathrm{warp}}=[d,\theta,\phi]$. These offsets are mapped into the token embedding space through a learnable projection function $f_{\mathrm{proj}}$:
@@
Appendix B Detailed Ablation Analysis
Effect of Discretization Bins. A key advantage of our model is its ability to flexibly discretize continuous control signals into an arbitrary number of bins, enabling a fine-grained trade-off between numerical precision and representation compactness. To evaluate this, we systematically vary the discretization resolution from 128 to 8192 bins (see Tab. 4). At low resolutions (e.g., 128 bins), the model fails to learn meaningful trajectories due to the inability of coarse quantization to accurately encode joint values. This results in erratic and unstable robot motions. As the bin count increases, control accuracy improves, with performance saturating around 2048 bins. At this point, the discretization error becomes negligible compared to diffusion noise, offering sufficient precision for smooth and stable control. However, further increasing the resolution to 4096 or 8192 bins yields only marginal gains or even slight degradation in performance. This is attributed to the larger token vocabulary and increased noise, which complicate optimization. Unlike conventional autoregressive VLA models limited to 256 discrete bins, our method supports significantly higher-resolution action representations, enabling superior control fidelity and precision.
@@ Appendix B Detailed Ablation Analysis
Effect of Action Dimension. Our model supports flexible action dimensionality when discretizing continuous control vectors. To assess the influence of this flexibility, we vary the number of action dimensions $d\in\{8,16,32\}$ during training and evaluation (see Tab. 5). We observe that setting $d$ to match or slightly exceed the maximum action dimensionality in the dataset (e.g., d=8d{=}8 in LIBERO) yields the highest average success rate. When $d$ is smaller than the true action dimension, the model cannot capture all control channels, leading to missing or corrupted actions during decoding. Conversely, using a value of $d$ significantly larger than necessary introduces numerous blank or zero-padded tokens into the sequence. These redundant tokens inject unnecessary noise into the diffusion process, degrading the sharpness and precision of predicted trajectories. Therefore, for optimal performance, $d$ should be chosen to tightly align with the dataset’s action space dimensionality, ensuring both representational completeness and robustness against embedding-induced noise.
@@
Effect of One-hot Smooth Factor. Prior to injecting noise into the one-hot action representation, we apply a one-hot smooth factor to scale the vector. This smoothing operation attenuates the sharp peaks of the one-hot distribution, thereby reducing the dominance of high-confidence logits and enabling smoother transitions within the action space during diffusion. As shown in Tab. 6, lower values of the smooth factor consistently improve success rates across all LIBERO subsets. Stronger attenuation encourages the model to better explore neighboring discrete actions and prevents overfitting to rigid one-hot boundaries. Conversely, higher smooth factor values retain excessively sharp logits, which hinder denoising stability and degrade generalization performance. These findings indicate that moderately suppressing one-hot amplitude prior to noise injection enhances the robustness of policy learning under diffusion-based action modeling.
@@
Effect of Action Chunk. We adopt an action chunking strategy, wherein the policy outputs a short sequence of future actions conditioned on the current observation. This approach has shown particular effectiveness in real-world imitation learning and VLA models (e.g., (Chi et al., 2023; Liu et al., 2024; Black et al., 2024)). By predicting multiple future actions simultaneously, the policy captures implicit temporal context, thereby mitigating non-Markovian ambiguities such as pauses, delayed responses, or lingering motions—without requiring additional historical inputs that may introduce causal confusion.
@@
Effect of Data Normalization. We compare two normalization techniques for processing continuous action vectors within our model. Both methods apply element-wise normalization based on statistical parameters (e.g., mean and standard deviation) computed independently for each action dimension across the entire dataset.
@@
Appendix C Hardware Details
We conducted real-world experiments using a Franka Emika Research 3 robotic arm, equipped with two Intel RealSense D435i cameras. One camera was mounted on the robot’s wrist, providing a dynamic view that changes with the motion of the robotic arm. The second camera was placed laterally in the workspace to capture a stable third-person perspective. We utilized Gello (Wu et al., 2024) to control the Franka arm and collect demonstration data.
@@ Appendix C Hardware Details
Appendix D More Details on LIBERO
To evaluate the generalization behavior of our model under diverse distribution shifts, we adopt LIBERO (Liu et al., 2023) as a comprehensive benchmark for robot manipulation. LIBERO provides 130 language-conditioned tasks grouped into four procedurally generated suites (spatial, object, goal, and libero-100), each targeting a different dimension of knowledge transfer, including spatial reasoning, object concepts, behavioral goals, and mixed declarative–procedural knowledge. These tasks are created through a scalable generation pipeline that samples human-activity-inspired templates, constructs scene layouts, and specifies PDDL-based goal predicates, resulting in highly diverse manipulation scenarios with controllable variations in object types, spatial layouts, and goal semantics.
Following the common practice in prior work (use libero-10 instead of libero-100) and consistent with the configurations adopted by most existing models (Black et al., 2024; Kim et al., 2024; Pertsch et al., 2025), we train a single unified policy across all LIBERO tasks, instead of training task-specific models. All training is performed using a single RTX PRO 6000 GPU over 30,000 steps. Each task provides 50 high-quality teleoperated demonstrations, enabling sample-efficient behavioral cloning and ensuring that performance differences primarily reflect the model’s ability to transfer and retain knowledge across tasks.
@@ Appendix D More Details on LIBERO
Diffusion Policy (Chi et al., 2023) 78.3 92.5 68.3 50.5 72.4
MDT (Reuss et al., 2024) 78.5 87.5 73.5 64.8 76.1
OpenVLA (Kim et al., 2024) 84.7 88.4 79.2 53.7 76.5
Octo (Mees et al., 2024) 78.9 85.7 84.6 51.1 75.1
Dita (Hou et al., 2025) 84.2 96.3 85.4 63.8 82.4
TraceVLA (Zheng et al., 2024) 84.6 85.2 75.1 54.1 74.8
SpatialVLA (Qu et al., 2025) 88.2 89.9 78.6 55.5 78.1
$\pi_{0}$ FAST (Pertsch et al., 2025) 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 94.2
$\pi_{0.5}$ (Intelligence et al., 2025) 95.4 98.4 97.2 89.6 95.2
$\mathcal{E}_{0}$ (ours) 97.2 99.4 95.0 92.2 96.0
Appendix E More Details on ManiSkill
To further assess the generalization capability and fine-grained manipulation performance of our model across diverse operational scenarios, we additionally conduct experiments on ManiSkill (Tao et al., 2024). ManiSkill is an advanced GPU–parallelized simulation and benchmarking framework for embodied AI, offering one of the most comprehensive suites of manipulation tasks to date. It covers 12 distinct categories, ranging from tabletop manipulation and dexterous hand control to room-scale mobile manipulation, bimanual coordination, and high-fidelity digital twin environments. Its highly efficient GPU-based simulation and rendering pipeline enables large-scale visual policy training on a single GPU with minimal memory overhead, making it a suitable platform for evaluating both generalization and visual robustness.
Following the experimental design used in RDT (Liu et al., 2024), we leverage the official demonstration dataset (5 tasks with a total of 5,000 trajectories) and retrain all baseline models under identical data conditions to ensure fair comparison. All training is performed on a single RTX PRO 6000 GPU for 30,000 steps, using the same evaluation protocol across models. This setup allows us to rigorously compare their representational capacity, generalization behavior, and fine-grained manipulation performance within a high-fidelity simulation environment.
@@ Appendix E More Details on ManiSkill
Diffusion Policy (Chi et al., 2023) 0.0 40.0 80.0 0.0 88.0 30.2
OpenVLA (Kim et al., 2024) 0.0 8.0 80.0 0.0 8.0 4.8
Octo (Mees et al., 2024) 0.0 0.0 0.0 0.0 0.0 0.0
RDT (Liu et al., 2024) 13.2 77.2 74.0 1.2 100.0 53.6
$\pi_{0}$ FAST (Pertsch et al., 2025) 0.0 80.0 52.0 0.0 92.0 44.8
$\pi_{0}$ (Black et al., 2024) 4.0 60.0 48.0 0.0 100.0 42.4
$\pi_{0.5}$ (Intelligence et al., 2025) 8.0 56.0 56.0 4.0 92.0 43.2
$\mathcal{E}_{0}$ (ours) 24.0 76.0 72.0 4.0 100.0 55.2
Appendix F More Details on VLABench
To further assess the robustness, semantic understanding, and generalization capability of our model across diverse language-conditioned manipulation scenarios, we additionally evaluate it on VLABench (Zhang et al., 2024). VLABench offers a broad and heterogeneous collection of tasks that span various object categories, spatial arrangements, and visually complex environments. Each task is paired with natural language instructions that require accurate semantic grounding, attribute-level understanding, and precise alignment between linguistic cues and visual observations. The benchmark’s extensive variability in object layouts, appearance conditions, and linguistic expressions makes it a rigorous platform for evaluating multimodal reasoning and visuomotor robustness.